A 6kg object travelling at 10m/s collides head on with a 4kg object travelling at 4m/s. Determine the velocity of 10 m/s. Assume that the collisions occur in an isolated island


1
Expert's answer
2021-06-14T17:27:14-0400
"v_{1f}=\frac{m_1-m_2}{m_1+m_2}v_{1i}+\frac{2m_2}{m_1+m_2}v_{2i}"

"v_{1f}=\frac{6-4}{6+4}10+2\frac{4}{6+4}(-4)=-1.2\frac{m}{s}"